40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 998     Entry time: Fri Sep 26 16:08:39 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: some progress 
There's been good progress in locking the last couple of nights. A lot of time was wasted before I found that
all the SUS{POS,PIT,YAW} damping gains on the SRM were set to 0.1 for some reason, which let it get rung up
just a bit during bang locking. After setting these gains to 0.5 (similar to PRM and BS), the initial lock
acquisition of DRMI+2ARMs (nospring) got much quicker. Then more time was wasted by sticky sliders on the
transmon QPD whitening gain, causing the Schmitt triggered HI/LO gain PD switch not to happen. This meant
that the arm power was not reported properly when the CARM offset was reduced, and so loop gain normalizations
were not working properly. After all this, by the end of the night last night, reduced the CARM offset such
that stored power in the arms was about half of the max. Should be able to get to full power with another
good night, and then back to springy locking.
ELOG V3.1.3-